OpenStudy (anonymous):

Can someone please help? This is the last question on my worksheet and I'm so stuck and completely brain dead. :( Find the exact value by using a half-angle identity. tan(7pi/8)

jimthompson5910 (jim_thompson5910):

hint: use the identity shown on this page http://www.mathwords.com/h/half_angle_identities.htm and notice how (1/2)*(7pi/4) = 7pi/8 so that means you'll be locating 7pi/4 on the unit circle

OpenStudy (anonymous):

Ah thank you for even looking at this! I've been here for like an hour hoping to get help...

jimthompson5910 (jim_thompson5910):

Sorry for the wait. Does the hint and/or the page help at all?

OpenStudy (anonymous):

So it's sqrt2/2?

jimthompson5910 (jim_thompson5910):

the unit circle is this http://www.regentsprep.org/regents/math/algtrig/ATT5/600px-Unit_circle_angles_svg.jpg

OpenStudy (anonymous):

Right and then I see that the one for 7pi/4 is sqrt2/2 and -sqrt2/2

jimthompson5910 (jim_thompson5910):

notice how tan(7pi/8) = -0.41421 when we use a calculator and make sure it's in radian mode so that means that the final answer must be negative

jimthompson5910 (jim_thompson5910):

that means if we use the identity \[\large \tan\left(\frac{x}{2}\right) = \pm \sqrt{\frac{1-\cos(x)}{1+\cos(x)}}\] then we have to use the negative form of it and we know to use this version below \[\large \tan\left(\frac{x}{2}\right) = -\sqrt{\frac{1-\cos(x)}{1+\cos(x)}}\] making sense so far?

OpenStudy (anonymous):

sure but then what was the point of the unit circle?

OpenStudy (anonymous):

what were we looking for if we're using this formula to solve the equation?

jimthompson5910 (jim_thompson5910):

well we use the unit circle to determine cos(x) in this case, x = 7pi/4 so you have the right idea to get cos(7pi/4) = sqrt(2)/2

OpenStudy (anonymous):

So now we plug sqrt2/2 in then...ok that makes sense

jimthompson5910 (jim_thompson5910):

so we'll have \[\large \tan\left(\frac{x}{2}\right) = -\sqrt{\frac{1-\cos(x)}{1+\cos(x)}}\] \[\large \tan\left(\frac{1}{2}*x\right) = -\sqrt{\frac{1-\cos(x)}{1+\cos(x)}}\] \[\large \tan\left(\frac{1}{2}*\frac{7\pi}{4}\right) = -\sqrt{\frac{1-\cos\left(\frac{7\pi}{4}\right)}{1+\cos\left(\frac{7\pi}{4}\right)}}\] \[\large \tan\left(\frac{1}{2}*\frac{7\pi}{4}\right) = -\sqrt{\frac{1-\frac{\sqrt{2}}{2}}{1+\frac{\sqrt{2}}{2}}}\] see how I'm getting all this so far?

OpenStudy (anonymous):